OpenStudy (anonymous):

6 0 urielshields if you substitute x= -1/2 on this equation 2x^2=7x+7, do you get the same answer on both side?

OpenStudy (anonymous):

2 (-1/2)^2 = 2*(1/4) = 1/2 7(-1/2) + 7 = -3 1/2 + 7 = 3 1/2

OpenStudy (anonymous):

no

OpenStudy (anonymous):

oh my bad I mean this equation: x= -1/2 on this equation 2x^2=7x+4